OpenStudy (anonymous):

i got 660 and i understand that the lcm is the smallest number that all the numbers will go into and with the 31 it will only go into 1

OpenStudy (amistre64):

the reply box is this one down here on the bottom right of the screen

OpenStudy (amistre64):

its got a post button, an equation button, and this area here to type in

OpenStudy (anonymous):

Im sorry

OpenStudy (amistre64):

its easier to have a conversation within the confines of a single post rather than have it strewn across multiple postings :)

OpenStudy (amistre64):

thats better :)

OpenStudy (anonymous):

im having problems finding the lcm and the gcf for the ages 10,31 and 44 can you please explain

OpenStudy (amistre64):

22 10 31 was our numbers right?

OpenStudy (amistre64):

44 10 and 31

OpenStudy (anonymous):

yes and can you explain as you go

OpenStudy (amistre64):

hmmm, one method is to break the numbers into prime factors does that make sense?

OpenStudy (amistre64):

22 = 2,11 10 = 2,5 31 = 1,31

OpenStudy (anonymous):

yes and i have the numbers broke down

OpenStudy (anonymous):

can you tell me if im right

OpenStudy (zarkon):

GCD(10,31,44 )=GCD(GCD(10,31),44)=GCD(1,44)=1

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

11*4= 44 2*5=10 31=31*1

OpenStudy (amistre64):

4 is not prime, break it again into 2*2 11*2*2 = 44 the others are good

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay now from there i need to find the LCM

OpenStudy (anonymous):

MY LEAST COMMON FACTORS ARE (10,31 and 44)

OpenStudy (amistre64):

2*2*11: two 2s, one 11 2*5 : one 2, one 5 1*31: one 1, one 31 since the 2s come up twice: one 2, and two 2s we take one that has the most: two 2s the rest of the primes occur once so they are just admitted in: 1*2*2*5*11*31

OpenStudy (anonymous):

so my lcm is 2

OpenStudy (amistre64):

this is a common way of explaining it; but I dont like it. I perfer my own method

OpenStudy (amistre64):

no, your lcm is what we get in the end of it, not halfway thru

OpenStudy (anonymous):

660

OpenStudy (anonymous):

after timing everything

OpenStudy (amistre64):

what I like to do is make fractions out of our numbers: \[\frac{1}{44},\frac{1}{10},\frac{1}{31}\] and add them up to equal some arbitrary number. \[\frac{1}{44}+\frac{1}{10}+\frac{1}{31}=N\] then I start clearing fractions and reducing as I go

OpenStudy (amistre64):

this to me makes alot more sense

OpenStudy (anonymous):

okay slow down

OpenStudy (anonymous):

sorry writing steps so i dont forget

OpenStudy (anonymous):

are you there